The Acme manufacturing company tests their products before shipping. In how many ways can they select both 4 Cliff-Jumpers from 10 and 3 Pole-Climbers from 7 that they have ready to ship?
Answer by stanbon(75887)

The Acme manufacturing company tests their products before shipping. In how many ways can they select both 4 Cliff-Jumpers from 10 and 3 Pole-Climbers from 7 that they have ready to ship?
---------
Select 4 from 10 : 10C4 = (10*9*8*7)/(1*2*3*4) = 210 ways
------
Select 3 from 7 : 7C3 = (7*6*5)/(1*2*3) = 35
-----
# of ways to do both: 210*35 = 7350
=======================================
Cheers,
Stan H.
==================